# The quest for the GRAph Level autoEncoder (GRALE)

### v1

@@ Abstract
Although graph-based learning has attracted a lot of attention, graph representation learning is still a challenging task whose resolution may impact key application fields such as chemistry or biology. To this end, we introduce GRALE, a novel graph autoencoder that encodes and decodes graphs of varying sizes into a shared embedding space. GRALE is trained using an Optimal Transport-inspired loss that compares the original and reconstructed graphs and leverages a differentiable node matching module, which is trained jointly with the encoder and decoder. The proposed attention-based architecture relies on Evoformer, the core component of AlphaFold, which we extend to support both graph encoding and decoding. We show, in numerical experiments on simulated and molecular data, that GRALE enables a highly general form of pre-training, applicable to a wide range of downstream tasks, from classification and regression to more complex tasks such as graph interpolation, editing, matching, and prediction.
@@ 1 Introduction
Scrutinizing the literature, we observe that most existing works on graph AutoEncoders provide node-level embeddings instead of one unique graph-level embedding (See Fig. 1, left). In the following, we refer to this class of models as Node-Level AutoEncoders. The most emblematic example is the celebrated VGAE model Kipf and Welling, (2016), where the encoder is a graph convolutional network (GCN) that returns node embeddings $z_{i}$, and the decoder reconstructs the adjacency matrix from a dot product $A_{i,j}=\sigma(\langle z_{i},z_{j}\rangle)$. Many extensions have been proposed for this model, such as adversarial regularization Pan et al., (2018) or masking Tu et al., (2023), and it has been shown to be efficient for many node-level tasks, such as node clustering Zhang et al., (2022, 2019); Wang et al., (2017) or node outlier detection Du et al., (2022). Node-level models can also be used for graph generation, given that one knows the size (number of nodes) of the graph to generate; this includes GVAE Kipf and Welling, (2016) but also GraphGAN Wang et al., (2018) and graph normalizing flows Liu et al., (2019). When a graph-level representation is needed, one can apply some pooling operation on the node embeddings, for instance $z=\sum z_{i}$. Yet, this strategy is not entirely satisfying, as it inevitably leads to information loss, and it becomes difficult to decode a graph back from this representation Vinyals et al., (2015); Winter et al., (2021).
@@ 2.1 Problem Statement
Permutation invariant encoder. The first challenge is a well-studied topic for which a variety of architectures have been proposed, most of which rely on message passing Zhou et al., (2020); Zhu et al., (2022); Xu et al., (2018) or attention-based Müller et al., (2023); Rampášek et al., (2022) mechanisms. In this work, we have chosen to use an attention-based architecture, the Evoformer module from AlphaFold Jumper et al., (2021). As the numerical experiments show it, this architecture is particularly powerful, enabling the encoding and updating of pairwise relationships between graph nodes. To maintain symmetry with the encoder, the decoder also uses a novel Evoformer Decoder module. Details are provided in Appendix C.
@@
$\mathcal{L}_{\text{PI}}(G,\hat{G})=\min_{P\in\sigma_{N}}\mathcal{L}_{\text{ALIGN}}(G,P[\hat{G}]),$ (1)
where $\sigma_{N}$ is the set of permutation matrices and $P[G]$ denote the application of permutation $P$ to graph $G$ i.e. $P[G]=(Ph,PF,PCP^{T})$. This can be seen as first, aligning the two graphs, and then computing a loss $\mathcal{L}_{\text{ALIGN}}$ between them. This is problematic because, for any nontrivial loss $\mathcal{L}_{\text{ALIGN}}$ such that $\mathcal{L}_{\text{ALIGN}}(x,y)=0\iff x=y$, the optimization problem in (1) is NP complete Fortin, (1996). We discuss how to avoid this pitfall in the next section.
@@ 2.2 Matching-free reconstruction loss
where $G=(h,F,C)$, $\hat{G}=(\hat{h},\hat{F},\hat{C})$, and $\ell_{h},\ell_{F},\ell_{C}$ are ground losses responsible for the correct prediction of node masking, node features and edge features respectively. Despite the relaxation, this loss still satisfies key properties as stated in Propositions 1 and 2 (proofs in Appendix F).
@@
3 Overview of GRALE architecture
@@ 3.3 Decoder $f$
The decoder architecture mirrors that of the encoder. First, a transformer encoder updates the graph representation $Z$, then a novel Evoformer decoder module reconstructs the output graph nodes and edges. This new module, detailed in Appendix C, is based on the original Evoformer module, augmented with cross-attention blocks to enable decoding. As before, $\hat{X}$ is obtained by applying a simple linear layer to the last layer hidden node representations.
@@ 3.4 Matcher $m$ and loss
$\hat{T}=m(X,\hat{X})$ (7)
To enforce that $\hat{T}\in\pi_{N}$, we decompose the matcher in two steps: 1) We construct an affinity matrix $K$ and $K_{i,j}=\texttt{Aff}(X_{i},\hat{X}_{j})$ where Aff is a learnable affinity function 2) We apply the Sinkhorn algorithm to project $K$ on $\pi_{N}$ in a differentiable manner i.e. $\hat{T}=\texttt{SINKHORN}(K)$ Eisenberger et al., (2022); Flamary et al., (2018); Genevay et al., (2018).
@@
$\mathcal{L}_{\text{GRALE}}(G,\phi,\psi,\theta)=\mathcal{L}_{\text{OT}}\bigg{(}G,\underbrace{f_{\text{graph}}^{\phi}\circ g_{\text{graph}}^{\psi}(G)}_{\hat{G}},\underbrace{m^{\theta}\Big{(}g_{\text{nodes}}^{\psi}(G),f_{\text{nodes}}^{\phi}\circ g_{\text{graph}}^{\psi}(G)\Big{)}}_{\hat{T}}\bigg{)}.$ (8)
The detailed implementation of the modules mentioned in this section is provided in Appendix C and the whole model is trained end-to-end with classic batch gradient descent as described in A.
@@ 4.1 Permutation Invariant Graph Variationnal AutoEncoder (PIGVAE)
$\mathcal{L}_{\text{PIGVAE}}(G,\hat{G},\hat{T})=\mathcal{L}_{\text{ALIGN}}(G,\hat{T}[\hat{G}])$ (9)
where $\mathcal{L}_{\text{ALIGN}}(G,\hat{G})=\sum_{i=1}^{N}\ell_{F}(F_{i},\hat{F}_{i})+\sum_{i,j=1}^{N}\ \ell_{C}(C_{i,j},\hat{C}_{i,j})$. This reveals that $\mathcal{L}_{\text{PIGVAE}}$ and $\mathcal{L}_{\text{OT}}$ can be seen as two different relaxations of the same underlying matching problem $\min_{P\in\sigma_{N}}\mathcal{L}_{\text{ALIGN}}(G,P[\hat{G}])$. We detail this relationship in Proposition 3. However, Proposition 4 highlights an important limitation of the relaxation chosen in PIGVAE as the loss can be zero without a perfect graph reconstruction. All proofs are provided in appendix F.
@@
$\mathcal{L}_{\text{PIGVAE}}(G,\hat{G},P)=\mathcal{L}_{\text{OT}}(G,\hat{G},P)=\mathcal{L}_{\text{ALIGN}}(G,P[\hat{G}])$ (10)
If we relax to $\hat{T}\in\pi_{N}$ we only have an inequality instead $\mathcal{L}_{\text{PIGVAE}}(G,\hat{G},\hat{T})\leq\mathcal{L}_{\text{OT}}(G,\hat{G},\hat{T})$.
@@ 5.1 Reconstruction performances and ablation studies
Model performance and impact of components.
First, we compare GRALE to the other graph-level AutoEncoders (PIGVAE Winter et al., (2021), GraphVAE Simonovsky and Komodakis, (2018)). Table 1 shows that GRALE outperforms the other models by a large margin. Next, we conduct ablation studies to assess the impact of the individual components of our model. For each component, we replace it with a baseline (detailed in E) and measure the effect on performance. Results are shown in Table 2. Overall, all new components contribute to performance improvements. We also report the variance of the reconstruction metrics over 5 training seeds, revealing that while the choice of loss function and featurizer has limited impact on average performance, it significantly improves training robustness.
@@ 5.2 Qualitative properties of GRALE
Capturing graph properties in the embedding.
We randomly sample 10,000 molecules from two molecular datasets (PUBCHEM 32 and QM9 Wu et al., (2018)), compute their GRALE embeddings, and apply dimension reduction techniques (PCA and T-SNE). The resulting 2D projections shown in Fig. 4 are colored by graph size, solubility (logP), and internal energy (available only for QM9). In all cases, the projections reveal a clear structure, illustrating GRALE’s ability to capture both geometrical and chemical graph-level properties.
Figure 4: 2D projection of graph embeddings learned by GRALE, each point is a graph and the color represents important properties such as graph size, solubility (logP), and internal energy (u0).
@@ 5.3 Quantitative performances of GRALE models on downstream tasks
Graph classification/Regression.
We use the graph representations obtained by GRALE as input for classification and regression tasks in the MoleculeNet benchmark Wu et al., (2018). We compare our method to several graph representation learning baselines, including graph AutoEncoders (PIGVAE Winter et al., (2021), VGAE Kipf and Welling, (2016)) and contrastive learning methods (Infograph Sun et al., (2019), Simgrace Xia et al., (2022)). For a fair comparison, all models are pre-trained on PUBCHEM 32, and the same predictive head is used for the downstream tasks. Detailed settings are provided in Appendix A. Overall, GRALE outperforms the other graph AutoEncoders and performs similarly, if not better, than the other baselines.
Graph prediction.
Graph Matching.
where we enforce that $T(G_{1},G_{2})\in\sigma_{N}$ by replacing the Sinkhorn algorithm with the Hungarian algorithm Edmonds and Karp, (1972) inside the matcher. Then, we use this (potentially suboptimal) matching to compute an upper bound of the edit distance. In Table 5, we compare this approach to more classical methods as implemented in Pygmtools Wang et al., (2024). Pygmtools is fully GPU compatible, which enables a fair comparison with the proposed approach in terms of time complexity. The matcher operates out of distribution since it is trained to match input/output graphs that are expected to be similar $G_{1}\approx G_{2}$ at convergence and here we sample random pairs of graphs. Despite this, our model outperforms the classical solvers as it returns better upper bounds of the edit distance and is typically orders of magnitude faster.
@@
SimGRACE $0.110\pm 0.003$ $0.025\pm 0.005$ $0.293\pm 0.020$ $0.534\pm 0.023$ $0.374\pm 0.008$ $\mathbf{0.745}\pm 0.072$ $\mathbf{0.866}\pm 0.050$
InfoGraph $0.122\pm 0.001$ $0.020\pm 0.001$ $\mathbf{0.255}\pm 0.016$ $\mathbf{0.495}\pm 0.013$ $0.297\pm 0.010$ $0.729\pm 0.053$ $\underline{0.845}\pm 0.046$
@@
GRALE $\mathbf{0.015}\pm 0.001$ $\mathbf{0.018}\pm 0.003$ $\underline{0.274}\pm 0.014$ $\underline{0.511}\pm 0.022$ $\mathbf{0.272}\pm 0.017$ $\underline{0.731}\pm 0.025$ $0.821\pm 0.051$
@@
Any2Graph 0.19 $1.22$ $\underline{2.13}$ $\underline{3.63}$ $84.44$ $43.77$ $29.85$ $16.25$
@@
GRALE + Finetuning $\underline{0.27}$ 0.45 2.61 2.25 90.04 88.87 35.27 53.22
Table 5: We report (avg±$\pm$±std) edit distance and compute time over 1000 pairs of test graphs. All solvers use the default Pygmtools parameters except Greedy A which is A with beam width one.
@@
Greedy A $20.41\pm 11.21$ $\mathbf{32.53}\pm 8.15$ $0.021\pm 0.016$ $0.110\pm 0.054$
A $\underline{19.66}\pm 11.01$ N.A. $3.487\pm 1.928$ >100
GRALE $\mathbf{08.92}\pm 05.61$ $\underline{32.77}\pm 11.67$ $\underline{0.005}\pm 0.001$ $\mathbf{0.008}\pm 0.002$
@@ A.1 Datasets
Dataset Graph Size: Avg±$\pm$ Std Graph Size: Max N samples
@@
Dataset Graph Size: Avg±$\pm$ Std Graph Size: Max N samples N samples original
@@
Dataset Graph Size: Avg±$\pm$ Std Graph Size: Max N samples
@@ A.2 Model and training hyperparameters
As detailed in the previous section, we train 3 variants of our models, respectively, on COLORING 20, PUBCHEM 32 and PUBCHEM 16. We report the hyperparameters used in Table 9. Note that to reduce the number of hyperparameters, we set the number of layers in all modules (evoformer encoder, transformer decoder, transformer encoder, evoformer decoder) to the same value $L$, idem for the number of attention heads $H$ and node/edge hidden dimensions. In all cases, we trained with ADAM Kingma and Ba, (2014), a warm-up phase, and cosine annealing. We also report the number of GPUs and the total time required to train the models with these parameters.

### v3

@@ Abstract
Although graph-based learning has attracted a lot of attention, graph representation learning is still a challenging task whose resolution may impact key application fields such as chemistry or biology. To this end, we introduce GRALE, a novel graph autoencoder that encodes and decodes graphs of varying sizes into a shared embedding space. GRALE is trained using an Optimal Transport-inspired loss that compares the original and reconstructed graphs and leverages a differentiable node matching module, which is trained jointly with the encoder and decoder. The proposed attention-based architecture relies on Evoformer, the core component of AlphaFold, which we extend to support both graph encoding and decoding. We show, in numerical experiments on simulated and molecular data, that GRALE enables a highly general form of pre-training, applicable to a wide range of downstream tasks, from classification and regression to more complex tasks such as graph interpolation, editing, matching, and prediction. [^1] Code available at https://github.com/KrzakalaPaul/GRALE
@@ 1 Introduction
Scrutinizing the literature, we observe that most existing works on graph AutoEncoders provide node-level embeddings instead of one unique graph-level embedding (See Fig. 1, left). In the following, we refer to this class of models as Node-Level AutoEncoders. The most emblematic example is the celebrated VGAE model Kipf and Welling, (2016), where the encoder is a graph convolutional network (GCN) that returns node embeddings $z_{i}$, and the decoder reconstructs the adjacency matrix from a dot product $A_{i,j}=\sigma(\langle z_{i},z_{j}\rangle)$. Many extensions have been proposed for this model, such as adversarial regularization Pan et al., (2018) or masking Tu et al., (2023), and it has been shown to be efficient for many node-level tasks, such as node clustering Zhang et al., (2022); Zhang et al., 2019b ; Wang et al., (2017) or node outlier detection Du et al., (2022). Node-level models can also be used for graph generation, given that one knows the size (number of nodes) of the graph to generate; this includes GVAE Kipf and Welling, (2016) but also GraphGAN Wang et al., (2018) and graph normalizing flows Liu et al., (2019). When a graph-level representation is needed, one can apply some pooling operation on the node embeddings, for instance $z=\sum z_{i}$. Yet, this strategy is not entirely satisfying, as it inevitably leads to information loss, and it becomes difficult to decode a graph back from this representation Vinyals et al., (2015); Winter et al., (2021).
@@ 2.1 Problem Statement
Permutation invariant encoder. The first challenge is a well-studied topic for which a variety of architectures have been proposed, most of which rely on message passing Zhou et al., (2020); Zhu et al., (2022); Xu et al., (2018) or attention-based Müller et al., (2023); Rampášek et al., (2022) mechanisms. In this work, we have chosen to use an attention-based architecture, the Evoformer module from AlphaFold Jumper et al., (2021). As the numerical experiments show it, this architecture is particularly powerful, enabling the encoding and updating of pairwise relationships between graph nodes. To maintain symmetry with the encoder, the decoder also uses a novel Evoformer Decoder module. Details are provided in Appendix D.
@@
[…]
where $\sigma_{N}$ is the set of permutation matrices and $P[G]$ denote the application of permutation $P$ to graph $G$ i.e. $P[G]=(Ph,PF,PCP^{T})$. This can be seen as first, aligning the two graphs, and then computing a loss $\mathcal{L}_{\text{ALIGN}}$ between them. This is problematic because, for any nontrivial loss the optimization problem in (1) is NP-complete Hartmanis, (1982). We discuss how to avoid this pitfall in the next section.
@@ 2.2 Matching-free reconstruction loss
where $G^=(h^,F^,C^)\hat{G}=(\hat{h},\hat{F},\hat{C}),$ and $\ell_{h},\ell_{F},\ell_{C}$ are ground losses responsible for the correct prediction of node masking, node features and edge features respectively. Despite the relaxation, this loss still satisfies key properties as stated in Propositions 1 and 2 (proofs in Appendix G).
@@
3 GRALE architecture
3.1 Overview
[…]
We next provide the details of this implementation.
@@ 3.4 Decoder $f$
The decoder architecture mirrors that of the encoder. First, a transformer encoder updates the graph representation $Z$, then a novel Evoformer decoder module reconstructs the output graph nodes and edges. This new module, detailed in Appendix D, is based on the original Evoformer module, augmented with cross-attention blocks to enable decoding. As before, $\hat{X}$ is obtained by applying a simple linear layer to the last layer hidden node representations.
@@ 3.5 Matcher $m$ and loss
[…]
To enforce that $\hat{T}\in\pi_{N}$, we decompose the matcher in two steps: 1) We construct a node affinity matrix $K_{i,j}=\texttt{Aff}(X_{i},\hat{X}_{j})$ where Aff is a learnable affinity function 2) We apply Sinkhorn projections (Algorithm 2) to project $K$ on $\pi_{N}$ in a differentiable manner i.e. $\hat{T}=\texttt{SINKHORN}(K)$ Eisenberger et al., (2022); Flamary et al., (2018); Genevay et al., (2018).
@@
[…]
The detailed implementation of the modules mentioned in this section is provided in Appendix D and the whole model is trained end-to-end with classic batch gradient descent as described in A.
Data: Graph $G$
# Encode
$Z,X=g_{\text{graph}}(G),g_{\text{nodes}}(G)$
# Decode
$\hat{G},\hat{X}=f_{\text{graph}}(Z),f_{\text{nodes}}(Z)$
# Match
[…]
# Loss
[…]
Result: $\ell$
Algorithm 1 GRALE Loss
Data: Node affinity $K$, Iterations $n_{\text{iter}}$
$u,v=\mathbf{0}$
For $i$ in $n_{\text{iter}}]$ do:
# Row normalization
$u=-(K+\mathbf{1}v^{T}).\operatorname{logsumexp}(\mathrm{dim}=2)$
# Column normalization
$v=-(K+u\mathbf{1}^{T}).\operatorname{logsumexp}(\mathrm{dim}=1)$
# Matching
$\hat{T}=\exp(K+u\mathbf{1}^{T}+\mathbf{1}v^{T})$
Result: $T$
Algorithm 2 Sinkhorn Projection
@@ 4.1 Permutation Invariant Graph Variationnal AutoEncoder (PIGVAE)
[…]
@@
[…]
@@ 5.1 Reconstruction performances and ablation studies
Model performance and ablations.
First, we compare GRALE to the other graph-level AutoEncoders (PIGVAE Winter et al., (2021), GraphVAE Simonovsky and Komodakis, (2018)). Table 1 shows that GRALE outperforms the other models by a large margin. Next, we conduct ablation studies to assess the impact of the individual components of our model. For each component, we replace it with a baseline (detailed in F) and measure the effect on performance. Results are shown in Table 2. Overall, all new components contribute to performance improvements. We also report the variance of the reconstruction metrics over 5 training seeds, revealing that while the choice of loss function and featurizer has limited impact on average performance, it significantly improves training robustness.
@@ 5.3 Quantitative performances of GRALE models on downstream tasks
SimGRACE $110\pm 0.003 0.0$$025\pm 0.005$ $0.293\pm 0.020$ $0.534\pm 0.023$ $0.374\pm 0.008$ $\underline{0.745}\pm 0.072$ $\mathbf{0.866}\pm 0.050$
InfoGraph $122\pm 0.001 0.0$$020\pm 0.001$ $\mathbf{0.255}\pm 0.016$ $\mathbf{0.495}\pm 0.013$ $0.297\pm 0.010$ $0.729\pm 0.053$ $0.845\pm 0.046$
GraphMAE $0.222\pm 0.016$ $0.247\pm 0.036$ $0.291\pm 0.012$ $0.527\pm 0.029$ $0.378\pm 0.028$ $\mathbf{0.773}\pm 0.036$ $\underline{0.857}\pm 0.039$
@@
GRALE $\mathbf{0.015}\pm 0.001$ $\mathbf{0.018}\pm 0.003$ $\underline{0.274}\pm 0.014$ $\underline{0.511}\pm 0.022$ $\mathbf{0.272}\pm 0.017$ $0.731\pm 0.025$ $0.821\pm 0.051$
Graph classification/Regression.
We use the graph representations obtained by GRALE as input for classification and regression tasks in the MoleculeNet benchmark Wu et al., (2018). We compare our method to several graph representation learning baselines, including graph AutoEncoders (PIGVAE Winter et al., (2021), VGAE Kipf and Welling, (2016), GraphMAE Hou et al., (2022)) and contrastive learning methods (Infograph Sun et al., (2019), Simgrace Xia et al., (2022)). For a fair comparison, all models are pre-trained on PUBCHEM 32, and the same predictive head is used for the downstream tasks. Detailed settings are provided in Appendix A. Overall, GRALE outperforms the other graph AutoEncoders and performs similarly, if not better, than the other baselines.
@@
Any2Graph 0.19 $1.22$ $\underline{2.61}$ $\underline{3.63}$ $84.44$ $43.77$ $29.85$ $16.25$
@@
GRALE + Finetuning $\underline{0.27}$ 0.45 2.13 2.25 90.04 88.87 35.27 53.22
Graph prediction.
[…]
@@
Greedy A $20.41\pm 11.21$ $\underline{32.53}\pm 08.15$ $0.021\pm 0.016$ $0.110\pm 0.054$
A $19.66\pm 11.01$ N.A. $3.487\pm 1.928$ >100
GRALE $\underline{08.92}\pm 05.61$ $32.77\pm 11.67$ $\underline{0.005}\pm 0.001$ $\mathbf{0.008}\pm 0.002$
GRALE + Finetuning $\mathbf{07.64}\pm 04.42$ $\mathbf{19.33}\pm 07.85$
Graph Matching.
where we enforce that $T(G_{1},G_{2})\in\sigma_{N}$ by replacing the Sinkhorn algorithm with the Hungarian algorithm Edmonds and Karp, (1972) inside the matcher. Then, we use this (potentially suboptimal) matching to compute an upper bound of the edit distance. In Table 5, we compare this approach to more classical methods as implemented in Pygmtools Wang et al., (2024). Pygmtools is fully GPU compatible, which enables a fair comparison with the proposed approach in terms of time complexity. The matcher operates out of distribution since it is trained to match input/output graphs that are expected to be similar $G_{1}\approx G_{2}$ at convergence and here we sample random pairs of graphs. To mitigate this, we fine-tune the matcher on random pairs of graphs from the training set. As shown in Table 5, the proposed approach yields better upper bounds of the edit distance than classical solvers while being orders of magnitude faster.
@@ A.1 Datasets
[…]
@@
[…]
@@
[…]
@@ A.2 Model and training hyperparameters
As detailed in the previous section, we train 3 variants of our models, respectively, on COLORING 20, PUBCHEM 32 and PUBCHEM 16. We report the hyperparameters used in Table 9. Note that to reduce the number of hyperparameters, we set the number of layers in all modules (evoformer encoder, transformer decoder, transformer encoder, evoformer decoder) to the same value $L$, idem for the number of attention heads $H$ and node/edge hidden dimensions. In all cases, we trained with ADAM Kingma and Ba, (2014), a warm-up phase, and cosine annealing. Finally, we report the training complexity for GRALE and its Graph-Level competitors in table 10.
@@
Table 10: Training complexity for the different graph-level AutoEncoders. We report the theoretical complexity and the empirical training time given the hyperparameters of Table 9.
Model COLORING PUBCHEM 16 PUBCHEM 32 Complexity
GraphVAE 40h 80h N.A. $O(N^{4})$
PIGVAE 8h 18h 95h $O(N^{3})$
GRALE 8h 20h 100h $O(N^{3})$